1. In the year 2010, the government decided to construct a thermal plant in a small town. Discuss its impact on equilibrium wage in the small town in the short and long-term. (5 Marks)


2. Kwabena earns $15 per hour for up to 40 hours of work each week. He is paid $30 per hour for every hour in excess of 40. Kwabena faces a 20 percent tax rate. Kwabena receives $80 from the government as part of the Livelihood Empowerment Against Poverty (LEAP) program which is tax-free. There are 110 (non-sleeping) hours in the week. Graph Kwabena's weekly budget line. (10 Marks)


3. Suppose that the supply curve for private schoolteachers is Ls = 20,000 + 350W, and the demand curve for such schoolteachers is Ld = 100,000 - 150W, where L = the number of teachers and W = the daily wage.

(a). Plot the supply and demand curves.

(b). What are the equilibrium wage and employment levels in this market?

(c). What will be the effect on labour demand and supply if the government imposes a minimum wage of 200 GHS. (10 Marks)


4. Akua gains utility from consumption C and leisure L. The most leisure she can consume in any given week is 110 hours. Her utility function is U (C, L) = C × L. Akua receives $660 each week from her great-grandmother—regardless of how much she works.

(a). What will be Akua's marginal rate of substitution.

(b). What will be Akua's reservation wage? (Explain in detail)


1
Expert's answer
2021-03-02T17:49:03-0500

1.In this case, the thermal plant can be a monopsony in the labor market. In the short term, in a monopsonic labor market, the marginal labor cost curve MRCL is higher than the labor supply curve S, so the marginal labor cost will be greater than the labor price. Maximizing profit, the monopsonist firm will change the amount of variable resource (labor) until it ensures equality of the marginal income from labor and the marginal cost of labor:

MRCL=MRPL

By equating the marginal cost of labor with the marginal return on labor, the monopsonist will choose the optimum point on the labor supply curve. A monopsony will employ LM units of labor compared to Lc in a market of perfect competition (LM< Lc).

At the same time, the company will pay a lower tariff rate than in a competitive market. In the long term condition of maximum profit in the long term is the proportionality limit products resources not their prices and marginal costs of firms on acquisition of resources.

Thus, the monopsony will always seek to underestimate the wages of potential employees.


2.Even if I worked 20 hours more per week, it means

"15*40+20*30-(15*40+20*30)*0.2+80=1040"

This is a weekly net income


3.

a)




b) Ls=Ld

20.000 + 350W=100,000 - 150W

500W=80 000

W=160

"L=20 000+350W=20 000+350\times160=76 000"

c)if the government imposes a minimum wage of 200 GHS.

W=200

"Ls=20 000+350W=20 000+350\times200=90 000"

"Ld=100 000-150W=100 000-150\times200=70 000"

if the minimum wage increases to 200, the supply will increase, the demand will fall


4.

a)The Marginal Rate of Substitution (MRS) is defined as the rate at which a consumer is ready to exchange a number of units good X for one more of good Y at the same level of utility.

"MRS=-\frac{\Delta y}{\Delta x}"

C=800

L=200

"MRS=-\frac{\Delta y}{\Delta x}=-\frac{800}{200}=-4"

b)reservation wage - the minimum wage that can stimulate employment (below which a person does not agree to work, for example, if the salary is the same or slightly higher than the unemployment benefit

The reservation wage is the MRS when not working at all. Thus, wRES = MRS at maximum leisure;

"wRes=\frac{ C}{ L} =\frac{660}{110}= 6"